OpenStudy (anonymous):

please help me, if x = -2, y = 1, and z = 5, 2xyz^3 + x^2 - z^2 is -521 or -471 ?????

OpenStudy (kira_yamato):

2xyz^3 + x^2 - z^2 = 2(-2)(1)(5)^3 + (-2)^2 - (5)^2 = (-4)(125) + 4 - 25 = -500 - 21 = -521

OpenStudy (anonymous):

thankyou @Kira_Yamato